acopeland27:

Solve 3.8 = p/3 - 7.2.

Extrinix:

\(3.8 = \dfrac{p}{3}-7.2\) Add 7.2 to both sides \(11=\dfrac{p}{3}\) Convert \(\dfrac{p}{3}\) to \(\dfrac{1}{3}p\) \(11=\dfrac{1}{3}p\) Multiply both sides by the reciprocal of \(\dfrac{1}{3}\), which is 3 \(33=p\)

acopeland27:

thankyou

Extrinix:

No problem ^^